OpenStudy (bbb911):

This is Geometry: Write an equation of a circle with center (-7,5) and radius=9

OpenStudy (anonymous):

Do you know the formula for the equation of a circle?

OpenStudy (thewayiam983):

its going to be 9^2=81=(x+7)sq +(y-5)sq i think

OpenStudy (bbb911):

yes thats correct thank you :D

OpenStudy (anonymous):

It's very easy: \[Circle: r^2 = (x-h)^2 + (y-k)^2\] Where r is the radius, h is the x-coordinate of the center, and k is the y-coordinate of the center.